OpenStudy (anonymous):

HELP Which correctly orders the numbers from least (smallest) to greatest (largest)? -0.875, 0.5, 3/5, 66.6% -0.5, 3/5, 0.875, 66.6% -3/5, 66.6%, 0.5, 0.875 -0.5, 3/5, 66.6%, 0.875

OpenStudy (anonymous):

is it the last one??

OpenStudy (unklerhaukus):

this is easy of you covert all the numbers to decimals \[\frac 35=\frac 6{10}=\frac{60}{100}=0.6\] \[66\%=\frac{66.6}{100}=0.666\]

OpenStudy (anonymous):

im kinda confused is it the 1st one or the last option?

OpenStudy (unklerhaukus):

first and last are both right

OpenStudy (unklerhaukus):

\[-0.875< 0.5< 0.6<0.666 \]also \[-0.5, 0.6, 0.666, 0.875\] are those \(-\)signs ment to be negative signs or are they just indicating the start of a new option, if they are ment to be negative signs then there are two correct answers but, now im thinking they are just ment to indicate the start of a new option, so only one of the option is going to be correct

OpenStudy (unklerhaukus):

so your were right the whole time, \[0.5< 0.6< 0.666< 0.875\]